OpenStudy (lucaz):

solution set xln(x) = ln(x^2)

OpenStudy (anonymous):

\[x\ln x-\ln x^2=0\]\[x\ln x-2 \ln x=0\]\[(x-2)\ln x=0\]zero product property

OpenStudy (lucaz):

ok, thanks!